Hydrochloric acid is usually purchases in concentr
Answer #1

Basis : 100 gms of hydrochloric acid solution.

Given the density = 1.19 g/ml, Volume of the solution = mass*density = 1.19*100 =119 ml =119/1000 =0.119L

Mass % of HCl is 37.4%, mass of HCl =37.4*100/100 gm= 37.4 gms

Molecular weight of HCl= 36.5, moles of HCl= mass/Molecular weight of HCl =37.4/36.5=1.024 moles

Molarity= moles/ L of solution = 1.024/0.119 = 8.61